Tom bought a desktop computer and a laptop computer. Before finance charges, the laptop cost
$350 more than the desktop. He paid for the computers using two different financing plans. For the desktop the interest rate was 6.5% per year, and for the laptop it was 9% per year. The total finance charges for one year were$388. How much did each computer cost before finance charges?

Answer by richwmiller(17219)

L-D=350
L=350+D
0.065*L+0.09*D=388
0.065*(350+D)+0.09*D=388
22.75+0.065D+0.09*D=388
0.155*D=365.25
D=2356.45 at 9%
L=350+D
L=2706.45 at 6.5%
2356.45161
check
0.065*2706.45161+0.09*2356.45161=388
175.919355+212.080645=388
388.0=388
ok